Suppose you place 0.0600 g of magnesium chips in a coffee-cup calorimeter and then add 100.0 mL of 1.00 M HCl. The reaction that occurs is Mg(s) + 2 HCl(aq) → H2(g) + MgCl2(aq) The temperature of the solution increases from 23.11 °C (296.26 K) to 25.81 °C (298.96 K). What is the enthalpy change for the reaction per mole of Mg? Assume a specific heat capacity of the solution is 4.20 J/g∙K and the density of the HCl solution is 1.00 g/mL. kJ/mol
Answer #1

The heat transferred in the reaction is calculated:

q = m * cp * ΔT = 100 * 4.20 * (25.81 - 23.11) = 1134 J

Mg moles are calculated:

n Mg = m / MM = 0.06 / 24.31 = 0.0025 mol

The reaction enthalpy change is calculated:

ΔH = - q / n = - 1,134 kJ / 0.0025 mol = - 453.6 kJ / mol

If you liked the answer, please rate it positively, you would help me a lot, thanks.